Rational Expressions and Applications:
Ms. Tseng, a high school mathematics teacher, gave a test on perimeter, area, and volume to her geometry classes. Working alone, it would take her 4 hr to grade the test. Her student teacher, Jonah Schmidt, would take 6 hr to grade the same test. How long would it take them to grade these tests if they work together?

Answer by jorel1380(3719)

Working alone, it would take Ms. Tseng 4 hours, so she does 1/4th the work in an hour. Likewise, Schmidt does 1/6th the work in an hour. Working together:
x/4+x/6=1 job.
-----------
x/4+x/6=1
3x+2x=12
5x=12
x=12/5
It would take 12/5ths hours for the two working together to finish grading the tests.